OpenStudy (anonymous):

check my math? x^5 = x^2 * x^3 x^6 = (x^-2)^3 x = x^-2x^3 x^-6 = (x^3)^2 1/x = x^2 * x^-3

TheSmartOne (thesmartone):

the second one isn't correct... and neither is the 4th one

TheSmartOne (thesmartone):

Everything else is correct :)

TheSmartOne (thesmartone):

If you switch the answers of the second and the fourth one then it will be correct...

OpenStudy (anonymous):

Thank you! c:

TheSmartOne (thesmartone):

x^6 = (x^2)^3 x^-6 = (x^-2)^3

TheSmartOne (thesmartone):

oops x^6=(x^3)^2

TheSmartOne (thesmartone):

it is the same thing but I switched the numbers around..

OpenStudy (anonymous):

ahh yeah. Thanks, you helped a lot!! :)

TheSmartOne (thesmartone):

My pleasure :)